OpenStudy (bahrom7893):

jmfranco PHYSICS TOPIC: Mass of earth Mass of moon Mass of sun Distance b/n earth and moon Distance b/n earth and sun Universal grav. constant Radius of earth

OpenStudy (bahrom7893):

okay so let me start on this!

OpenStudy (bahrom7893):

guys id really appreciate if someone just googled those values and pasted them here, please? im helpin too many ppl at the same time

OpenStudy (bahrom7893):

\[M_e = 5.9742 × 10^{24} kg\] \[M_m = 7.36 \times 10^{22} kg\] \[M_s = 1.98892 \times 10^{30}\]

OpenStudy (anonymous):

ok

OpenStudy (bahrom7893):

\[D_{em} = 3.840 \times 10^5 km\] = \[D_{em} = 3.840 \times 10^8 m\]

OpenStudy (bahrom7893):

hey when is this due?

OpenStudy (anonymous):

tomoro

OpenStudy (bahrom7893):

gimme ur email, typing stuff out is a nightmare, i will write it out take a pic and send it to u

OpenStudy (anonymous):

ok its jackie.franco12@yahoo.com that would be great

OpenStudy (anonymous):

thanksss

OpenStudy (bahrom7893):

email?

OpenStudy (bahrom7893):

k workin on it

OpenStudy (anonymous):

ya ok thanks

OpenStudy (bahrom7893):

k done with part one for the moon, will take a pic an send

OpenStudy (bahrom7893):

k done with everything off to bed, emailed u btw, nite